# Double white dwarf binary population in MOCCA star clusters

### v2

@@ 4.2 Wide binaries
In this section, we report our results involving wide binaries that have been obtained by filtering our mocca data according to 3.5.2. We split our results into clusters with initial binary fraction of 10% and 95%. As we discussed in Sect. 4.1, our limit in magnitude causes high mass WDs to become undetectable at later snapshots.
4.2.1 10% Binary fraction
@@
Figure 4 shows the distribution of projected separations plotted against mass of both binary components for 10% initial binary fraction at 2, 5 and 9 Gyr. For all snapshots, the binaries from mocca are not as wide as the observed binaries. This is because when we use 10% binary fraction, we have a cut-off at 100 au for initial semi-major axis. This is done under the assumption that the soft components of the binary populations are entirely gone and the remaining binaries are hard binaries. However, recent studies have shown that having nearly 100% initial binaries with most being soft is a better assumption. (e.g. Leigh et al. 2015; Belloni et al. 2019)
@@
It is therefore unlikely that the binaries will evolve into larger separations. Since the absence of wide systems is seen in all three datasets, including the standalone evolution set, this cannot be due to dynamical interactions. Looking at the mass distribution we can see that the masses from our simulations, as expected, depend on time. Over time, lower mass WDs binaries are being formed as expected so for later times the distribution of masses gets lower, in addition to this, the cooling process causes high mass WDs to be undetectable at later times. We find that at 5 Gyr, our distribution aligns fairly well with the data from Heintz et al. (2022). At 2 Gyr the masses are biased towards higher masses while for 9 Gyr they are biased towards lower masses due to the cooling of WDs.
4.2.2 95% binary fraction
@@
When using 95% binary fraction we allow wider binaries to be formed initially. This causes our separation distribution to be more shifted towards wider binaries. This is shown in Fig. 5. All 3 datasets from mocca are very similar for all snapshots but we will point out the small differences. At 2 Gyr we have a tail extending towards wider binaries but our distribution is still biased towards lower separations than the observations. The masses do not align very well at this snapshot, they are too high and the observed WDs have noticeably lower masses. Escapers and isolated binaries are very similar in their separation distribution while in-cluster binaries have slightly lower separations which could be due to disruptions in dynamical interactions.
@@
At 5 Gyr the separation distribution is very similar to the 2 Gyr snapshot. However, the masses agree much better with observations. Due to more time passing, lower mass WDs can form. In addition, the cooling sequence removes high mass WDs from our sample. This causes a good agreement with observations in mass. At 9 Gyr, the separation distribution does not change in any significant way. However, the cooling process has now made the higher mass WDs undetectable and we are left with a small subset of WDs with mass around $0.6$ M.
Similarly to the close binaries, we can see that over time the the number of binaries increases in the escapers and standalone evolution datasets while decreasing in the cluster dataset.
@@
Figure 4: Projected separation plotted against mass for wide binaries with 10% initial binary fraction. See Fig. 1 for an explanation of panels, colours and markers.
@@
Figure 5: Projected separation plotted against mass for wide binaries with 95% initial binary fraction. See Fig. 1 for an explanation of panels, colours and markers.
@@ 4.2.3 Heintz age distribution
Figure 6: Projected semi-major axis plotted against mass of the WDs for our set with an age distribution similar to Heintz et al. (2022). Panel (a) for 10% and (b) for 95%. See Fig. 1 for an explanation of colours and markers.
@@
We can see in the previous plots that there is a time dependence in the mass distribution, both due to lower mass WDs being formed and higher mass WDs not being detectable due to the cooling process. This plot shows a sample of our DWDs with the same age distribution as in Heintz et al. (2022). This is explained in Sect. 3.5.2. We can see that by doing this we indeed get a fairly good mass distribution, our masses are only slightly smaller but there is a bias and a peak at around $0.9$ M. This peak is connected to bse and in more detail to the envelope type. bse handles both convective and radiative envelopes but, as seen in this figure, the switch between may be too abrupt to be realistic. Nevertheless, if we would disregard this peak, our mass distribution agrees very well with observations.
@@
However, this does not solve the problem of too low separations. This is most likely due to observational bias and WD binaries with lower separations are not observed or included in this observational survey. While the observations have a slight bell curve shape with a peak around 800 au, the mocca data extends to much lower separations and we are only looking at the far end tail. Due to this, we add an additional cut and only include binaries with projected separation above 500 au. This can be seen in Fig. 7 where our agreement between separation is much better. With these filtering procedures we can reproduce the observations in a very good way, in both projected separation and mass for all 3 of our simulated datasets.
@@
Figure 7: Projected separation plotted against mass of the WDs in the binary. In this figure we added an additional cut-off at 500 au to remove the tail of lower separation binaries found in the observational data. We see good agreement in both mass and separation. See Fig. 1 for an explanation of colours and markers.
5 Summary & Conclusions
@@
In this paper, we have taken a large number of DWDs from a large number of mocca simulations and compared them to observed DWDs. This proved to be difficult due to strong observational biases. However, with some filtering and cuts on the mocca data, we managed to get good agreement in separation for close binaries and for both separation and masses for wide binaries. One important point about our data that should not be forgotten is that we are incorporating multiple stellar populations in our initial model. In these models, we have a first population this is tidally filling and not too concentrated and a second population that is significantly more concentrated than the first population. Thus our initial setup differs greatly from the traditional approach of having one dense initial population. In these new setups, we find a much larger number of escapers compared to simulations using the traditional one-population setup. We saw that many clusters lose 30-40% of their mass in just the first few Myr due to objects escaping the cluster induced by stellar evolution mass loss. Due to the early escape of these objects they are also mostly (or in many cases, completely) unaffected by dynamics before their escape. We can see this by the similarities in results between the ’standalone evolution’ and ’escapers’ data sets. This means that DWDs born in the field and those born inside GCs are, on average, similar.
@@
Our GC initial conditions give us both tidally filling and underfilling clusters. Since tidally filling clusters will lose more mass in the early stages of the evolution, the number of escaping DWDs is expected to be different for these two different types as well. We find that from tidally filling clusters, there are approximately 2 times more DWD escapers at 2 Gyr and approximately 1.6 times more DWD escapers at 9 Gyr compared to tidally underfilling clusters. However, we do not find any significant differences in the statistical distribution of separations of DWDs that have escaped from tidally filling and underfilling clusters.
@@
A significant finding is related to the number of DWDs in each set; we found that for most clusters and snapshots, the number of DWDs in the escaping dataset is far larger than the other two sets (in-cluster and standalone binaries). This is clearly shown in figure 8 where we plot the ratio between the number of escaping DWDs and standalone DWDs. For 95% binary fraction, we have slightly more escaping DWDs at 1 Gyr than standalone DWDs, however, at later times this ratio grows and we see a faster increase in the number of escaping DWDs. At 12 Gyr, we have approximately 3 times more escaping DWDs than standalone ones. For 10% binary fraction, we start with significantly less escaping DWDs. At 3 Gyr, we have approximately the same number and at later times we have more escaping than standalone. At 12 Gyr, we have almost 4 times more escaping DWDs than standalone. This indicates that even though many binaries escape early, they are still influenced by dynamics in a way that causes them to be more likely to form WDs within a Hubble time. This is true even for less dense clusters and clusters with a low binary fraction. For populations 1 and 2 separately, we can see that the big difference in this ratio comes from population 1, especially for 95% binary fraction at early times.
@@
Figure 8: Evolution of ratio between number of escaping DWDs and standalone DWDs for every Gyr between 1 and 12 Gyr. We have 6 groups in this figure; solid lines show both populations together, dashed lines show population 1 only and dotted lines show population 2 only.
@@
When looking at the whole range of semi-major axis (see Fig. 1 and LABEL:fig:aMass_095) we can see a gap in the in-cluster population between $\sim$0.5 and $\sim$25 au. For the escape and standalone populations, we do not see this clear gap. This can be seen more clearly in Fig. 9 where we plot a histogram of the distribution of semi-major axis for the three data sets from mocca at 2 Gyr (panel a) and 9 Gyr (panel b) for 95% binary fraction. We also split populations 1 and 2 into separate groups in order to see the effect of the two populations. In the plot we clearly see the gap in the in-cluster set for both population 1 (dark-red) and population 2 (orange) at both times. Comparing the two populations at 2 Gyr, for small separations we have approximately 55% higher fraction of population 2 binaries than population 1 binaries. However, at larger separations, we have approximately 25% higher fraction of population 1 binaries than population 2 binaries. At 9 Gyr the two populations have evened out and the differences are much smaller.
@@
This gap is non-existent for the escapers and standalone data sets. Instead, there is a sharp increase by $\sim$1 au. This big peak is mostly visible for escapers at 2 Gyr but is there for both data sets at both snapshots. This very big population of escaping binaries at $\sim$1 AU seems to indicate that this gap is partly due to escaping binaries, however, the reason for these particular binaries to escape is still uncertain. The fact that we do see a similar peak, however, not as large, for the standalone runs seems to indicate that we have some bias in our initial datasets to create WD binaries with these separations. Combining this with a large mass loss in the first few Myr can lead to features like this. However, this paper aimed to compare our results to observations and finding a more conclusive answer to this gap is beyond the scope of this paper. We can also note that, from these plots, there are no significant differences in populations 1 and 2 for the escapers and standalone sets. Even though there is no statistical difference in the distribution of separation between populations 1 and 2, there is a clear number difference. We have almost 100 times more escapers from population 1 than 2, at both 2 and 9 Gyr. This causes population 1 to provide the large bulk of escaping binaries when analysing the data as a whole.
@@
Figure 9: Histogram of the semi-major axis distribution for the three mocca datasets with population 1 and 2 split into different groups from a snapshot at 9 Gyr for 95% binary fraction.
Our main conclusions can be summarised as:
@@
  • Observations of DWD systems are biased and limited which make direct comparisons impossible without filtering mocca data.
@@
  • Using 95% binary fraction with initial semi-major axis distribution given by Belloni et al. (2017) allows us to form wider binaries that agree better with observations (see Fig. 5)
  • Using a similar age distribution as in Heintz et al. (2022) and adding a lower limit to the separation causes our 95% binary fraction dataset to agree very well with observations in both separation and masses (see Fig. 7)
  • The masses of WDs in wide binaries agree well with observations at 5 Gyr, but are too high at 2 Gyr and too low at 9 Gyr. This is true for both 10% and 95% binary fraction.
  • For close binaries, our data agrees well with observed binaries in separation. In particular for escapers and isolated binaries but the in-cluster binaries also agree fairly well (see Fig. 2 and 3).
  • The very low masses of ELM WDs are not able to be reproduced with bse. Uncertainties in binary evolution related to magnetic braking, mass stability criteria and mass transfer modelling makes it impossible to form these extremely low mass WDs.
@@
  • The cooling sequence of WDs and the magnitude limits of observations indicate the mostly young WDs are being observed. This is confirmed by the age estimates from Heintz et al. (2022) where a majority of binaries have an age less than 3 Gyr.
  • There is room for future analysis on these subjects, in particular the gap that we found in the separations of DWDs. The focus on this paper was on comparisons to observations and we were not able to find a concrete explanation to this gap. Also, with future observational missions there will be more data available. This would lead to better comparisons to simulated data, especially since most observational data is constrained to either close or wide binaries.
Acknowledgments
LH, GM, AH, AA, GW were supported by the Polish National Science Center (NCN) through the grant UMO-2016/23/B/ST9/02732. DB acknowledges financial support from FONDECYT grant number 3220167.
@@ Acknowledgments
Input and output data for the globular cluster simulations carried out in this paper will be shared on request to the corresponding author.  [^3] https://zenodo.org/records/10865904

### v3

@@ 4.2 Wide binaries
In this section we report our results involving wide binaries that were obtained by filtering our MOCCA data according to 3.5.2. We split our results into clusters with initial binary fractions of 10% and 95%. As we discussed in Sect. 4.1, our limit in magnitude causes high mass WDs to become undetectable at later snapshots.
4.2.1 Binary fraction of 10%
@@
Figure 4 shows the distribution of projected separations plotted against mass of both binary components for a 10% initial binary fraction at 2, 5, and 9 Gyr. For all snapshots, the binaries from MOCCA are not as wide as the observed binaries. This occurs when we use a 10% binary fraction; we have a cutoff at 100 au for initial semi-major axis. This is done under the assumption that the soft components of the binary populations are entirely gone and the remaining binaries are hard binaries. However, recent studies have shown that having nearly 100% initial binaries with most being soft is a better assumption (e.g., Leigh et al. 2015; Belloni et al. 2019).
@@
It is therefore unlikely that the binaries will evolve into larger separations. Since the absence of wide systems is seen in all three datasets, including the standalone evolution set, this cannot be due to dynamical interactions. Looking at the mass distribution we can see, as expected, that the masses from our simulations depend on time. Over time, lower mass WDs binaries are formed as expected, so for later times the distribution of masses becomes lower, in addition to this, the cooling process causes high mass WDs to be undetectable at later times. We find that at 5 Gyr our distribution aligns fairly well with the data from Heintz et al. (2022). At 2 Gyr the masses are biased toward higher masses, while at 9 Gyr they are biased toward lower masses due to the cooling of WDs.
4.2.2 Binary fraction of 95%
@@
When using a 95% binary fraction, we allow wider binaries to be formed initially. This causes our separation distribution to be more shifted toward wider binaries. This is shown in Fig. 5. All three datasets from MOCCA are very similar for all snapshots, but we point out the small differences. At 2 Gyr we have a tail extending toward wider binaries, but our distribution is still biased toward lower separations than the observations. The masses do not align very well at this snapshot: they are too high and the observed WDs have noticeably lower masses. Escapers and isolated binaries are very similar in their separation distribution, while in-cluster binaries have slightly smaller separations, which could be due to disruptions in dynamical interactions.
@@
At 5 Gyr the separation distribution is very similar to the 2 Gyr snapshot. However, the masses agree much better with observations. Due to more time passing, lower mass WDs can form. In addition, the cooling sequence removes high mass WDs from our sample. This causes a good agreement with observations in mass. At 9 Gyr, the separation distribution does not change in any significant way. However, the cooling process has now made the higher mass WDs undetectable and we are left with a small subset of WDs with mass around $0.6$ M. Similarly to the close binaries, we can see that over time the number of binaries increases in the escapers and standalone evolution datasets, while the number decreases in the cluster dataset.
Refer to captionRefer to captionRefer to caption
@@
Figure 4: Projected separation plotted against mass for wide binaries with a 10% initial binary fraction. See Fig. 1 for an explanation of panels, colors, and markers.
Refer to captionRefer to captionRefer to caption
@@
Figure 5: Projected separation plotted against mass for wide binaries with a 95% initial binary fraction. See Fig. 1 for an explanation of panels, colors, and markers.
@@ 4.2.3 Heintz age distribution
Figure 6: Projected semi-major axis plotted against mass of the WDs for our set with an age distribution similar to Heintz et al. (2022). See Fig. 1 for an explanation of colors and markers.
@@
We can see in the previous plots that there is a time dependence in the mass distribution, which is the result of lower mass WDs being formed and of higher mass WDs not being detectable due to the cooling process. Figure 6 shows a sample of our DWDs with the same age distribution as in Heintz et al. (2022). This is explained in Sect. 3.5.2. We can see that by doing this we indeed get a fairly good mass distribution; our masses are only slightly lower, but there is a bias and a peak at around $0.9$ M. This peak is connected to BSE, and is in more detail than the envelope type. BSE handles both convective and radiative envelopes, but, as seen in this figure, the switch between them may be too abrupt to be realistic. Nevertheless, if we disregard this peak, our mass distribution agrees very well with the observations.
@@
However, this does not solve the problem of too small separations. This is most likely due to observational bias and WD binaries with smaller separations are not observed or included in this observational survey. While the observations have a slight bell curve shape with a peak around 800 au, the MOCCA data extends to much smaller separations, and we are only looking at the far end tail. Due to this, we added an additional cut and only include binaries with projected separation above 500 au. This can be seen in Fig. 7 where our agreement between separations is much better. With these filtering procedures we can reproduce the observations in a very good way, in both projected separation and mass for all three of our simulated datasets.
@@
Figure 7: Projected separation plotted against mass of the WDs in the binary. In this figure we added an additional cutoff at 500 au to remove the tail of smaller separation binaries found in the observational data. We see good agreement in both mass and separation. See Fig. 1 for an explanation of colors and markers.
5 Summary and conclusions
@@
In this paper we took a large number of DWDs from 197 MOCCA simulations and compared them to observed DWDs. This proved to be difficult due to strong observational biases. However, with some filtering and cuts on the MOCCA data, we managed to get good agreement in separation for close binaries and for both separation and masses for wide binaries. One important point about our data is that we incorporated multiple stellar populations in our initial model. In these models we have a first population that is tidally filling and not too concentrated and a second population that is significantly more concentrated than the first population. Thus, our initial setup differs greatly from the traditional approach of having one dense initial population. In these new setups we find a much larger number of escapers compared to simulations using the traditional one-population setup. We saw that many clusters lose 30-40% of their mass in just the first few Myr due to objects escaping the cluster induced by stellar evolution mass loss. Due to the early escape of these objects, they are also mostly unaffected (or in many cases, completely unaffected) by dynamics before their escape. We could see this by the similarities in results between the standalone evolution dataset and escapers dataset. This means that DWDs born in the field and those born inside GCs and later ejected are, on average, similar.
@@
Our GC initial conditions gave us both tidally filling and underfilling clusters. Since tidally filling clusters lose more mass in the early stages of evolution, the number of escaping DWDs was expected to be different for these two different types as well. We find that from tidally filling clusters, there are approximately two times more DWD escapers at 2 Gyr and approximately 1.6 times more DWD escapers at 9 Gyr compared to tidally underfilling clusters. However, we do not find any significant differences in the statistical distribution of separations of DWDs that have escaped from tidally filling and underfilling clusters.
@@
A significant finding is related to the number of DWDs in each set. We found that for most clusters and snapshots, the number of DWDs in the escaping dataset is far larger than the other two sets (in-cluster and standalone binaries). This is clearly shown in Fig. 8 where we plot the ratio of the number of escaping DWDs to standalone DWDs. For a 95% binary fraction, we have slightly more escaping DWDs at 1 Gyr than standalone DWDs; however, at later times this ratio grows and we see a faster increase in the number of escaping DWDs. At 12 Gyr, there are approximately three times more escaping DWDs than standalone ones. For the 10% binary fraction, we start with significantly fewer escaping DWDs. At 3 Gyr we have approximately the same number, and at later times we have more escaping than standalone. At 12 Gyr, we have almost four times more escaping DWDs than standalone. This indicates that even though many binaries escape early, they are still influenced by dynamics in a way that causes them to be more likely to form WDs within a Hubble time. This is true even for less dense clusters and clusters with a low binary fraction. For populations 1 and 2 separately, we can see that the big difference in this ratio comes from population 1, especially for the 95% binary fraction at early times.
@@
Figure 8: Evolution of ratio of escaping DWDs to standalone DWDs for every Gyr between 1 and 12 Gyr. We have six groups in this figure. The solid lines show both populations together; the dashed lines show population 1 only; and the dotted lines show population 2 only.
@@
When looking at the whole range of semi-major axes (see Fig. 1), we can see a gap in the in-cluster population between $\sim$0.5 and $\sim$25 au. For the escape and standalone populations, we do not see this clear gap. This can be seen more clearly in Fig. 9 where we plot a histogram of the distribution of semi-major axes for the three datasets from MOCCA at 2 Gyr (panel a) and 9 Gyr (panel b) for the 95% binary fraction. We also split populations 1 and 2 into separate groups in order to see the effect of the two populations. In the plot we clearly see the gap in the in-cluster set for both population 1 (dark red) and population 2 (orange) at both times. Comparing the two populations at 2 Gyr, for small separations we have an approximately 55% higher fraction of population 2 binaries than population 1 binaries. However, at larger separations, we have an approximately 25% higher fraction of population 1 binaries than population 2 binaries. At 9 Gyr the two populations have evened out and the differences are much smaller.
@@
This gap is nonexistent for the escapers and standalone datasets. Instead, there is a sharp increase of $\sim$1 au. This big peak is mostly visible for escapers at 2 Gyr, but is there for both datasets at both snapshots. This very big population of escaping binaries at $\sim$1 AU seems to indicate that this gap is partly due to escaping binaries; however, the reason why these particular binaries escape is still uncertain. The fact that we see a similar peak, although not as large, for the standalone runs seems to indicate that we have some bias in our initial datasets to create WD binaries with these separations. Combining this with a large mass loss in the first few Myr can lead to features like this. However, this our aim was to compare our results to observations, and finding a more conclusive answer to this gap is beyond the scope of this paper. We can also note that, from these plots, there are no significant differences in populations 1 and 2 for the escapers and standalone sets. Even though there is no statistical difference in the distribution of separation between populations 1 and 2, there is a clear number difference. We have almost 100 times more escapers from population 1 than 2, at both 2 and 9 Gyr. This causes population 1 to provide the large bulk of escaping binaries when analyzing the data as a whole.
@@
Figure 9: Histogram of the semi-major axis distribution for the three MOCCA datasets. Populations 1 and 2 are split into different groups from a snapshot at 9 Gyr for the 95% binary fraction.
Our main conclusions can be summarized as follows:
@@
  • Observations of DWD systems are biased and limited, which make direct comparisons impossible without filtering MOCCA data.
@@
  • Using a 95% binary fraction with the initial semi-major axis distribution given by Belloni et al. (2017) allows us to form wider binaries that agree better with observations (see Fig. 5).
  • Using an age distribution similar to that used in Heintz et al. (2022) and adding a lower limit to the separation causes our 95% binary fraction dataset to agree very well with observations in both separation and mass (see Fig. 7).
  • The masses of WDs in wide binaries agree well with observations at 5 Gyr, but are too high at 2 Gyr and too low at 9 Gyr. This is true for both the 10% and 95% binary fractions.
  • For close binaries, our data agrees well with observed binaries in separation, in particular for escapers and isolated binaries; the in-cluster binaries also agree fairly well (see Figs. 2 and 3).
  • The very low masses of ELM WDs are not able to be reproduced with BSE. Uncertainties in binary evolution related to magnetic braking, mass stability criteria, and mass transfer modeling make it impossible to form these extremely low mass WDs.
@@
  • The cooling sequence of WDs and the magnitude limits of observations indicate the mostly young WDs are being observed. This is confirmed by the age estimates from Heintz et al. (2022), where a majority of binaries have an age younger than 3 Gyr. We are hoping that this conclusion holds for open younger clusters, which are also losing members that are slightly affected by dynamics.
  • There is room for future analysis on these subjects, in particular the gap that we found in the separations of DWDs. The focus on this paper was on comparisons to observations, and we were not able to find a concrete explanation to this gap. In addition, with future observational missions there will be more data available. This will lead to better comparisons with simulated data, especially since most observational data is constrained to either close or wide binaries.
Acknowledgements.
We would like to thank the reviewer for providing comments and suggestions that helped to improve the quality of the manuscript. LH, MG, AH, AA, GW were supported by the Polish National Science Center (NCN) through the grant UMO-2021/41/B/ST9/01191. DB acknowledges financial support from FONDECYT grant number 3220167. AA acknowledges support for this paper from project No. 2021/43/P/ST9/03167 co-funded by the Polish National Science Center (NCN) and the European Union Framework Programme for Research and Innovation Horizon 2020 under the Marie Skłodowska-Curie grant agreement No. 945339. For the purpose of Open Access, the authors have applied for a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
@@ Acknowledgements.
Input and output data for the globular cluster simulations carried out in this paper is available at https://zenodo.org/records/10865904